# Can Global XAI Methods Reveal Injected Bias in LLMs? SHAP vs Rule Extraction vs RuleSHAP

### v1

@@ Abstract
Generative AI systems can help spread information but also misinformation and biases, potentially undermining the UN Sustainable Development Goals (SDGs). Explainable AI (XAI) aims to reveal the inner workings of AI systems and expose misbehaviours or biases. However, current XAI tools, built for simpler models, struggle to handle the non-numerical nature of large language models (LLMs). This paper examines the effectiveness of global XAI methods, such as rule-extraction algorithms and SHAP, in detecting bias in LLMs. To do so, we first show a text-to-ordinal mapping strategy to convert non-numerical inputs/outputs into numerical features, enabling these tools to identify (some) misinformation-related biases in LLM-generated content. Then, we inject non-linear biases of varying complexity (univariate, conjunctive, and non-convex) into widespread LLMs like ChatGPT and Llama via system instructions, using global XAI methods to detect them. This way, we found that RuleFit struggles with conjunctive and non-convex biases, while SHAP can approximate conjunctive biases but cannot express them as actionable rules. Hence, we introduce RuleSHAP, a global rule extraction algorithm combining SHAP and RuleFit to detect more non-univariate biases, improving injected bias detection over RuleFit by +94% (MRR@1) on average.
@@ 1 Introduction
The United Nations’ Sustainable Development Goals (SDGs) address global challenges such as poverty (SDG 1), education (SDG 4), clean water (SDG 6), inequality (SDG 10), and climate change (SDG 13), each with specific targets for a sustainable future [1]. Meanwhile, AI has the potential to both accelerate and decelerate progress toward these goals [2]. For instance, [3] found limitations in the depth and reliability of generative AI for climate communication (SDG 13).
While large language models (LLMs) are widely used for knowledge discovery and education, their potential to spread biased or inaccurate information poses risks [4]. For instance, an LLM might dismiss a promising technology, such as the use of micro-algae for CO2 absorption, as ineffective, deterring users or investors from pursuing such solutions solely based on an LLM’s misrepresentation. Hence, to ensure AI contributes positively to the SDGs, it is critical to improve the quality of generated information in widespread LLM-powered applications. This will help align, among other things, AI technologies with the global sustainability agenda.
The complexity of LLM training pipelines [5, 6], coupled with the lack of publicly available training data [7, 8], makes it hard to analyse these models and detect biases arising from skewed or polluted datasets. Moreover, biases can be introduced (whether unintentionally or maliciously) through system instructions or other simple methods, especially by third-party users deploying ChatGPT or similar services in their applications. Consequently, identifying bias in LLMs requires a generic, black-box-compatible approach, motivating the use of explainable AI (XAI).
However, existing XAI tools, originally designed for simpler AI systems, struggle to scale to state-of-the-art LLMs [9], limiting our ability to detect biases in contexts like the UN SDGs. For example, SHAP [10] uses Shapley values to identify influential input features. While robust at the global scale, it relies on numerical inputs and ordinal outputs, unlike the vocabulary-based input/output space of LLMs. In contrast, RuleFit [11] uses gradient boosting [12] to extract rule-based relationships and LASSO [13] to prune them. Although RuleFit supports both categorical and ordinal spaces, it lacks SHAP’s theoretical guarantees and still struggles with the input/output space of LLMs.
To investigate global bias in widespread LLMs (i.e., ChatGPT, Llama) and detect biases, we first introduce text-to-ordinal mapping, a strategy to abstract the input and output spaces of LLMs into numerical, ordinal features, enabling the application of methods like SHAP and RuleFit. Inspired by the recent LLM-as-a-judge paradigm [14], our ordinal features are specifically crafted to reflect LLMs’ implicit judgments (i.e., opinions) on text properties linked to cognitive biases. These include phenomena such as framing effects, information overload, and oversimplification, key factors in the dissemination of misinformation on social networks [15, 16].
Hence, we present an automated assessment of global XAI algorithms, evaluating their ability to detect that kind of opinion-based biases in LLMs. Our assessment is based on the manual injection of biases via system instructions (preliminary prompts that shape how a LLM responds). We introduce three levels of bias complexity (univariate threshold, conjunctive multi-threshold, and non-convex/disconnected) to test the algorithms’ performance under varied conditions. We then evaluate how many of these rules are detected by state-of-the-art XAI methods with public implementation.
While SHAP often identifies the injected biases, it cannot express them as actionable, symbolic rules. Instead, methods like RuleFit produce too many rules and often fail to pinpoint non-univariate biases precisely. To address these shortcomings, we propose RuleSHAP, a novel algorithm that integrates SHAP into RuleFit by refining rule extraction and LASSO regression. Our results show that RuleSHAP outperforms all baselines in detecting non-linear biases injected into state-of-the-art LLMs. The replication package is available at: https://github.com/Francesco-Sovrano/RuleSHAP.
@@ 2 Related Work
Few model-agnostic XAI methods can extract global, non-linear rules from ordinal inputs/outputs (cf. Section 5), and all of them rely on tree surrogates. To our knowledge, RuleSHAP is the first to integrate global SHAP values to guide model-agnostic rule generation in tree surrogates. Indeed, while [17, 18, 19] produce global explanations, they do not support symbolic rule derivation.
Recent studies explore biases in LLMs using mixed methods [3, 20, 21] or focusing on demographic fairness [22]. For instance, [23] assess ChatGPT’s SDG literacy, and [24] examine human–LLM alignment challenges. Both emphasize attitudinal alignment, contrasting with our focus on detecting and ranking injected biases. [3], instead, highlight generative AI’s limitations in climate communication (SDG 13) only. In contrast, our framework automates data collection and bias analysis, supporting extensibility to new LLMs and abstraction schemes.
The most closely related work is probably [25], which also uses LLM-as-a-judge for bias detection but focuses on gender bias, validating that these metrics align with human judgments on response bias but without resorting to explicit rule extraction. Other works address cognitive or domain-specific biases without adopting global rule-extraction for LLMs. [26] examine anchoring and framing biases, proposing self-help interventions, while [27] study LLMs as text-quality judges. In the medical domain, [28] introduce BiasMedQA to quantify clinical biases.
3 Text-to-Ordinal Mapping for Bias Detection
@@
Preliminaries on global XAI. Two model-agnostic XAI approaches motivate our work: (i) perturbation-basedand (ii) rule-extractionmethods. Perturbation-based methods, e.g., LIME [29] and SHAP [10], analyse model responses to perturbed inputs $x^{\prime}$ generated by modifying subsets of features in $x\in\mathbb{R}^{d}$. SHAP estimates the feature importance $\rho_{i}$ to the output so as to satisfy properties like additivity: $f(x)=\sum_{i=1}^{d}\rho_{i}$. Rule-extraction methods, such as RuleFit [11], derive instead interpretable rules $R=\{r_{1},...,r_{k}\}$ from data points $X$ using surrogate models $f_{s}$ (e.g., decision trees), approximating the original model $f$ via $f_{s}(X)\approx f(X)$. Both approaches generally assume: (i) numerical or discrete input features suitable for perturbation; (ii) ordinal/categorical outputs.
Focus on bias detection within explanations. We seek to reveal biases that may affect how widespread general-purpose LLMs explain critical SDG themes. By focusing on topics where accurate, unbiased information is crucial, we can detect high-level biases that play a central role in the dissemination of misinformation [15, 16], such as framing effects (emphasizing certain attributes of an issue while downplaying others), information overload (producing overly long or complex outputs), and oversimplification (providing superficial explanations for complex matters).
@@
Formally, let $\mathcal{T}=\{\tau_{1},\tau_{2},...,\tau_{m}\}$ be a set of topics (i.e., textual phrases), each tied to an SDG-related challenge or solution (e.g., “micro-algae for CO2 absorption”). Let $\mathcal{X}={\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}}\subset\Sigma^{*}$ denote a set of prompt texts, each being a request for an explanation of a topic $\tau_{i}\in\mathcal{T}$ generated by a function $q\colon\mathcal{T}\to\mathcal{X}$ such that $\mathbf{x}_{i}=q(\tau_{i})$, where $q(\tau_{i})=$Explain $\tau_{i}$}” is the simplest choice. Then, $f:\mathcal{X}\to\Sigma^{*}$ models an LLM producing a textual explanation $\mathbf{y}_{i}=f(q(\tau_{i}))$ for each topic $\tau_{i}$.
@@
Abstraction setting. While a naive approach could rely on a vocabulary-based encoding that treats text as categorical tokens, it would fail to capture corpus-level biases. To address this limitation in the context of bias detection, we propose extracting domain-relevant properties from both the topics and their corresponding explanations. We define two mappings, $\phi_{x}\colon\mathcal{T}\to\mathbb{R}^{d}$ and $\phi_{y}\colon\Sigma^{*}\to\mathbb{R}^{m}$, that transform textual data into numerical (ordinal) representations: $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$ and $\mathbf{v}_{i}=\phi_{y}(\mathbf{y}_{i})\in\mathbb{R}^{m}$. Once these mappings are applied, the textual prompts and responses ($\mathbf{x}_{i}$, $\mathbf{y}_{i}$) are no longer used in the subsequent XAI pipeline. Under this framework, perturbation-based or rule-extraction XAI methods operate directly on the numeric vectors $\{\mathbf{u}_{i},\mathbf{v}_{i}\}$, thereby facilitating bias detection and interpretability in a LLM.
@@
Main assumptions and important workaround. The mapping process assumes that certain conceptual properties are learned in an AI’s latent space to model text generation. For instance, topics like war, disease, and famine often carry a negative bias in human writing, and we expect an LLM trained on such texts to share this bias. So, if bias relates to these properties, text-to-ordinal mapping can reveal the nature of this relationship. Moreover, the framework also assumes that topics with identical $\mathbf{u}$-vectors ($\mathbf{u}_{i}=\mathbf{u}_{j}$) yield similar output vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, an assumption that may fail due to LLM non-determinism. To address this, a sufficiently large $\mathcal{T}$ is needed to ensure statistical power and detectable correlations between $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$. These correlations can be estimated via Pearson or Spearman methods for linear or monotonic biases, or via distance correlation [30] for non-linear biases (as in our case). Thus, any global XAI insight from text-to-ordinal mappings should be supported by correlation evidence certifying abstraction validity.
@@
Input abstractions. Our analysis focuses on topics $\tau_{i}\in\mathcal{T}$ relevant to the UN SDGs. Recall that each topic $\tau_{i}\in\mathcal{T}$ is mapped to an ordinal feature vector $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$. In a standard LLM scenario, we do not have direct access to the training corpus. Hence, we propose using the LLM itself as $\phi_{x}(\cdot)$: as a judge providing informed opinions about a topic. In this way, an XAI algorithm can summarise how LLM’s explicit opinions on topic properties reflect underlying (opinion-based) biases. To assess whether different types of topics elicit skewed explanations, we define a set of 11 input properties mapped by $\phi_{x}(\cdot)$. These capture how the LLM perceives factors such as a topic’s prevalence in public discourse (commonality), emotional tone (positivity, negativity, neutrality), and breadth across fields (interdisciplinary nature). Additional properties include conceptual complexity, controversy, and regional and temporal variation. For a detailed breakdown of all properties, the specific prompts used to elicit LLM opinions, and other hyperparameters, see Appendix A. Specifically, LLMs are asked to rate, on a scale from 1 (low) to 5 (high), how a topic is typically portrayed on the web, e.g., whether it is common, positive, controversial, etc. Thus, for each topic $\tau_{i}$, we obtain a vector $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$ ($d=11$), where each component is an ordinal score on a 1–5 Likert scale.
@@
Output abstractions. Given a set of LLM-generated textual responses $\mathbf{y}_{i}=f(q(\tau_{i}))$, we define 7 output features to quantify relevant properties of each explanation $\mathbf{y}_{i}$. All these features are mapped to numerical scores by $\phi_{y}(\cdot)$. Thus, for each output $\mathbf{y}_{i}$, we construct a vector $\mathbf{v}_{i}=\phi_{y}(\mathbf{y}_{i})\in\mathbb{R}^{m}$ ($m=7$) capturing these properties. The first four features are proxy metrics commonly associated with the biases under study: (i) Explanation length (characters count) reflects information overload (if high) or oversimplification (if low) [31, 32]. (ii) Subjectivity and (iii) sentiment scores capture subjectivity and polarity, highlighting framing effects [33]. (iv) Readability indices, such as the Gunning fog index [34], may suggest instead information overload or oversimplification. The corresponding scoring functions $\phi_{y}(\cdot)$ are: a well-known readability index (Gunning fog index) yielding continuous scores [34]; character count as the measure of explanation length; publicly available sentiment/subjectivity estimators to generate sentiment/subjectivity scores (for more details see Appendix D). The remaining three features directly estimate (v) information overload, (vi) oversimplification, and (vii) framing effects. Inspired by [25], for these features we follow the strategy used with $\phi_{x}(\cdot)$, asking the LLM to rate (from 1 to 5) how much an explanation can produce a given bias (see Appendix B). This approach helps cross-validate whether our text-to-ordinal mapping captures bias-related signals. Indeed, if there is a bias, the direct estimates are expected to correlate with the four proxy metrics.

### v2

@@ Abstract
Large language models (LLMs) can amplify misinformation, undermining societal goals like the UN SDGs. We study three documented drivers of misinformation (valence framing, information overload, and oversimplification) which are often shaped by one’s default beliefs. Building on evidence that LLMs encode such defaults (e.g., “joy is positive,” “math is complex”) and can act as “bags of heuristics,” we ask: can general belief-driven heuristics behind misinformative behaviour be recovered from LLMs as clear rules? A key obstacle is that global rule-extraction methods in explainable AI (XAI) are built for numerical inputs/outputs, not text. We address this by eliciting global LLM beliefs and mapping them to numerical scores via statistically reliable abstractions, thereby enabling off-the-shelf global XAI to detect belief-related heuristics in LLMs. To obtain ground truth, we hard-code bias-inducing nonlinear heuristics of increasing complexity (univariate, conjunctive, nonconvex) into popular LLMs (ChatGPT and Llama) via system instructions. This way, we find that RuleFit under-detects non-univariate biases, while global SHAP better approximates conjunctive ones but does not yield actionable rules. To bridge this gap, we propose RuleSHAP, a rule-extraction algorithm that couples global SHAP-value aggregations with rule induction to better capture non-univariate bias, improving heuristics detection over RuleFit by +94% (MRR@1) on average. Our results provide a practical pathway for revealing belief-driven biases in LLMs.
@@ 1 Introduction
The United Nations’ Sustainable Development Goals (SDGs) address challenges such as poverty (SDG 1), clean water (SDG 6), and climate action (SDG 13). AI systems, and large language models (LLMs) in particular, are increasingly used for knowledge access, with the potential to accelerate or hinder progress towards these societal goals [2]. Empirically, their reliability in high-stakes domains remains uneven; e.g., Bulian et al. [3] document limitations in generative AI for climate communication. Beyond factual errors, LLMs may also propagate biased or misleading narratives that influence public understanding of SDG-relevant topics [4].
We focus on the latter, using bias to mean “a preference or inclination that inhibits impartial judgment,” and study three well-documented, misinformation-related mechanisms: valence framing, information overload, and oversimplification [5, 6]. These behaviours often stem from default or global beliefs [7, 8]. For instance, default polarities (e.g., “disease is negative”) can induce valence framing; perceived complexity (e.g., “quantum theory is hard”) can trigger overload or oversimplification. Notably, LLMs have been shown to encode global beliefs [9] and to operate via “bags of heuristics” [10]. This motivates our question: can belief-driven heuristics behind these behaviours (e.g., “if a topic is controversial, oversimplify it”) be recovered from LLMs as symbolic rules?
Global explainable AI (XAI) methods aim to extract such rule sets and are natural candidates, but most are optimized for numerical or tabular data and struggle with textual data [11]. For instance, SHAP [12] uses game theory to identify influential input features yet faces combinatorial explosions over token permutations in vocabulary-sized spaces [13]. RuleFit [14], instead, builds rules via gradient boosting and LASSO pruning, but lacks SHAP’s theoretical guarantees and still falters on the input/output spaces of LLMs.
Inspired by evidence that even arithmetic prompts trigger abstraction heuristics (e.g., based on parity or range) in LLMs [10], we present a statistically grounded abstraction pipeline that maps LLM beliefs (which are, by definition, self-expressed) and bias-estimation proxy metrics into numerical scores. This enables global XAI tools to detect belief-driven, bias-inducing behaviours in LLMs. To address LLM non-determinism, we collect thousands of LLM inputs and outputs with sample sizes determined through a priori power analysis. We then validate that the extracted rules and related abstractions capture genuine LLM behavioural regularities using correlation certificates.
We evaluate state-of-the-art global XAI methods (with publicly available implementations) on popular LLMs (ChatGPT, Llama) and on SDG-related topics. To establish ground truth, we hard-code 14 bias rules into these LLMs via three system instructions. To evaluate performance under varying nonlinearities, we analyse three rule types of increasing complexity: univariate, conjunctive, and non-convex. Thence, we assess how many rules are detected by the XAI methods.
While global SHAP flags injected bias, it cannot express it as rules. RuleFit instead overproduces rules and misses many non-univariate effects. We thus introduce RuleSHAP, which refines RuleFit’s rule generation via global SHAP-value aggregations. In controlled injections, RuleSHAP outperforms all baselines ($p\!\ll\!.05$), also cutting rules overproduction by 10.9% on average. Notably, rule-detection accuracy falls as rule complexity increases, and even our injected heuristics (simpler than many real-world cases) challenge state-of-the-art XAI, exposing critical limitations for future work to address. Our replication package is available at: https://github.com/Francesco-Sovrano/RuleSHAP.
@@ 2 Related Work
Global, model-agnostic XAI methods that recover non-linear rules for generic inputs/outputs are rare (§5); most almost uniformly rely on tree surrogates. To our knowledge, RuleSHAP is the first to use global SHAP values to steer model-agnostic rule induction within a tree surrogate.
Prior SHAP-based global explainers [15, 16, 17] provide importance attributions but neither symbolic rules nor explanations of LLMs. For autoregressive text generation (e.g., LLMs), SHAP faces a combinatorial explosion over token permutations, rendering exact computation intractable and undermining approximations [13, 18]. To mitigate this, Amara et al. [19] propose local concept-level abstractions in place of isolated tokens, an idea that aligns with the input/output abstractions we adopt.
Beyond SHAP‐based methods, mechanistic interpretability seeks to explain an LLM’s behaviour by finding neural circuits (neuron subsets driving outputs). But such work often fails to assign symbolic meaning or to separate local versus global behaviours [20]. An exception is Nikankin et al. [10], who show that for basic arithmetic LLMs rely on a “bag of heuristics”: identifiable neurons implement simple rules, and ablating them suppresses the associated behaviour.
Conversely, most LLM-bias studies use mixed-methods and centre on human–LLM alignment rather than extracting and ranking biases; e.g., Raman et al. [21] and Bulian et al. [3] test ChatGPT’s SDG literacy, Wu et al. [22] survey alignment challenges, while Kumar et al. [23] and Koo et al. [24] study LLMs as text-quality judges. In contrast, our framework automates data collection and bias analysis while remaining easily extensible to new LLMs and abstraction schemes.
3 LLM’s Belief Abstraction Pipeline for Rule-Extraction
@@
Preliminaries: Global XAI. Two model-agnostic XAI paradigms underpin our work: perturbation-based techniques and rule-extraction methods. Perturbation-based methods, e.g., LIME [25] and SHAP [12], analyse model responses to perturbed inputs $x^{\prime}$ generated by modifying subsets of features in $x\!\in\!\mathbb{R}^{d}$. SHAP estimates the feature importance $\rho_{i}$ to the output, satisfying properties like additivity: $f(x)\!=\!\sum_{i=1}^{d}\rho_{i}$. Rule-extraction methods, such as RuleFit [14], derive instead interpretable rules $R\!=\!\{r_{1},...,r_{k}\}$ from a dataset $X$ using surrogate models $f_{s}$ (e.g., decision trees), approximating the original model $f$ via $f_{s}(X)\!\approx\!f(X)$. Both approaches generally assume: (i) input features suitable for perturbation; (ii) outputs with an inherent ordering (e.g., logits, log-odds, probability distributions).
Data selection. In this paper, we use global XAI to surface high-level LLM heuristics that can propagate misinformation [5] by triggering three cognitive biases: framing effects (selective valence framing of attributes), information overload (overly long or complex outputs), and oversimplification (superficial explanations of complex issues). Existing datasets on cognitive bias in LLMs are confined to medicine [26, 4] and software engineering [27], and, critically, do not expose the underlying bias-inducing heuristics themselves. This lack of ground-truth rules prevents validation of explanations produced by global XAI; without them, the faithfulness of such explanations cannot be assessed.
@@
As no suitable dataset exists for our experiments, we focus on detecting heuristics that bias LLM-generated explanations of SDG-related topics, due to their global impact (e.g., misinformation on poverty or gender equality can have UN-wide implications) and the ease of generating related prompts. Formally, let $\mathcal{T}=\{\tau_{1},\tau_{2},...,\tau_{m}\}$ be a set of topics (i.e., textual phrases), each tied to an SDG-related challenge or solution (e.g., “micro-algae for CO2 absorption”). Let $\mathcal{X}={\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}}\subset\Sigma^{*}$ denote a set of prompt texts, each being a request for an explanation of a topic $\tau_{i}\in\mathcal{T}$ generated by a function $q\colon\mathcal{T}\to\mathcal{X}$ such that $\mathbf{x}_{i}=q(\tau_{i})$, where $q(\tau_{i})=$Explain $\tau_{i}$}” is the simplest choice. Then, $f:\mathcal{X}\to\Sigma^{*}$ models an LLM producing a textual explanation $\mathbf{y}_{i}=f(q(\tau_{i}))$ for each topic $\tau_{i}$.
@@
Abstraction setting. Motivated by evidence that LLMs encode data-induced global beliefs that often align with human judgments [28, 9], we extract abstract belief-related properties from topics (e.g., $\tau_{i}$ is common”) and proxy metrics for bias estimation from LLM-generated explanations (e.g., $\mathbf{y}_{i}$ is unreadable”). Specifically, we define two mappings, $\phi_{x}\colon\mathcal{T}\to\mathbb{R}^{d}$ and $\phi_{y}\colon\Sigma^{*}\to\mathbb{R}^{m}$, that transform textual data into numerical (ordered) representations: $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$ and $\mathbf{v}_{i}=\phi_{y}(\mathbf{y}_{i})\in\mathbb{R}^{m}$. Once these mappings are applied, the textual prompts and responses ($\mathbf{x}_{i}$, $\mathbf{y}_{i}$) are no longer used in the subsequent XAI pipeline. Under this framework, perturbation-based or rule-extraction XAI methods operate directly on the numeric vectors $\{\mathbf{u}_{i},\mathbf{v}_{i}\}$.
@@
Main assumption and important workaround. The framework above assumes that topics with identical $\mathbf{u}$-vectors ($\mathbf{u}_{i}=\mathbf{u}_{j}$) yield similar output vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, an assumption that may fail due to LLM non-determinism when scoring. To address this, a sufficiently large $\mathcal{T}$ is needed to ensure statistical power and detectable correlations between $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$. These correlations are estimated via Pearson or Spearman methods for linear or monotonic biases, or via distance correlation [29] for non-linear biases (as in our case). Thus, any global XAI insight from text-to-ordered mappings should be (and will be) supported by correlation evidence certifying abstraction validity.
@@
Input abstractions. We analyse SDG-related topics $\tau_{i}\in\mathcal{T}$. Each topic $\tau_{i}$ is mapped to an ordered feature vector $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$. Because $\mathbf{u}_{i}$ encodes the LLM’s belief about $\tau_{i}$, it follows that $\phi_{x}(\cdot)$ is, by definition, the LLM itself. To assess whether different types of topics elicit skewed explanations, we define a set of 11 input properties mapped by $\phi_{x}(\cdot)$. These capture how the LLM perceives factors such as a topic’s prevalence in public discourse (commonality), emotional tone (positivity, negativity, neutrality), and breadth across fields (interdisciplinary nature). Additional properties include conceptual complexity, controversy, and regional and temporal variation. For a detailed breakdown of all properties, the specific prompts used to elicit LLM opinions, and other hyperparameters, see Appendix A. Specifically, LLMs are asked to rate, on a scale from 1 (low) to 5 (high), how they think a topic is typically portrayed on the web, e.g., whether it is common, positive, controversial, etc. Thus, for each topic $\tau_{i}$, we obtain a vector $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$ ($d=11$), where each component is an ordered score on a 1–5 Likert scale.
@@
Output abstractions. Given a set of LLM-generated textual responses $\mathbf{y}_{i}=f(q(\tau_{i}))$, we define 7 output features to quantify relevant properties of each explanation $\mathbf{y}_{i}$. All these features are mapped to numerical scores by $\phi_{y}(\cdot)$. Thus, for each output $\mathbf{y}_{i}$, we construct a vector $\mathbf{v}_{i}=\phi_{y}(\mathbf{y}_{i})\in\mathbb{R}^{m}$ ($m=7$) capturing these properties. The first four features are proxy metrics commonly associated with the biases under study: (i) Explanation length (characters count) reflects information overload (if high) or oversimplification (if low) [30, 31]. (ii) Subjectivity and (iii) sentiment scores capture subjectivity and polarity, highlighting framing effects [32]. (iv) Readability indices, such as the Gunning fog index [33], may suggest instead information overload or oversimplification. The corresponding scoring functions $\phi_{y}(\cdot)$ are: a well-known readability index (Gunning fog index) yielding continuous scores [33], character count as the measure of explanation length, and publicly available sentiment/subjectivity estimators to generate sentiment/subjectivity scores (for more details see Appendix B.1).
Given evidence that larger LLMs (e.g., GPT-4) can align with human judgments [34], the remaining three features (only) are direct estimates of: (v) overload, (vi) oversimplification, and (vii) framing. Specifically, under an LLM-as-a-judge protocol [35], each explanation is rated (1–5) on how strongly it exhibits each bias (Appendix B.2). If the biases are real, these scores will correlate with the LLM-independent proxies, mitigating judge-hallucination concerns.